OpenStudy (anonymous):

Medal Solve 2(x + 1) = 2x + 2. All real numbers No solution 0 1

OpenStudy (anonymous):

a

OpenStudy (anonymous):

all real numbers

OpenStudy (nethercreep333):

Let's solve your equation step-by-step. 2(x+1)=2x+2 Step 1: Simplify both sides of the equation. 2(x+1)=2x+2 Simplify: (Show steps) 2x+2=2x+2 Step 2: Subtract 2x from both sides. 2x+2−2x=2x+2−2x 2=2 Step 3: Subtract 2 from both sides. 2−2=2−2 0=0 Answer: All real numbers are solutions.